# Magnetic Field Amplification and Particle Acceleration in Weakly Magnetized Trans-relativistic Electron-ion Shocks

### v1

@@ 3.2 Cosmic Ray Driven Instabilities
This corresponds to the simulation frame magnetic field energy density divided by the upstream kinetic energy density. Note that this is a frame-dependent quantity. The strength of the near upstream Weibel field seen at $(x-x_{\mathrm{sh}})/d_{\mathrm{i}}=[200,500]$, in the dark green and turquoise curves is $\varepsilon_{B}\sim 2\times 10^{-4}$. The shock transition region of the $\sigma=10^{-4.5}$ case, as seen from the $\varepsilon_{B}$ profile, is slightly broader, compared to the $\sigma=10^{-4}$ case. This is due to the larger-scale structures generated during the strongly nonlinear stage of the Weibel shock evolution (see Appendix A and Groselj2024). For the $\sigma=10^{-3.5}$ case in the orange curve, we see some spikes peaking at $\varepsilon_{B}\sim 10^{-3}$ at $(x-x_{\mathrm{sh}})/d_{\mathrm{i}}=[200,300]$, which is what we observed in Figure 3(c). The upstream magnetic field of the $\sigma=10^{-3}$ case is much stronger at $\varepsilon_{B}\sim 10^{-2}$. The high-current Bell modes generated in the $\sigma=10^{-3.5}$ case saturate an order of magnitude below the low-current Bell modes of the $\sigma=10^{-3}$ case, but still remain locally larger than the Weibel modes, which dominate at lower magnetizations. In all cases, $\varepsilon_{B}$ approaches the far upstream value;
@@
In this Subsection, we analyzed the cosmic ray currents and the plasma instabilities driven by cosmic rays in the upstream region. By considering the Weibel instability and both low-current and high-current Bell instability, we were able to elucidate the $\sigma$-dependence on the shock structure seen in Subsection 3.1. For additional details regarding the saturation of the high-current Bell modes, see Appendix B.
3.3 Particle Acceleration
@@
For electrons, the time evolution of the nonthermal spectra has a clear dependence on $\sigma$. For the low-$\sigma$ cases (panels (f,h)), the nonthermal tail has a relatively steady normalization, and its upper cutoff keeps growing. On the other hand, for the Bell-dominated case with $\sigma=10^{-3}$ (panel (b)), the normalization of the tail gets suppressed after $\omega_{\mathrm{pi}}t\sim 4000$. This is because the Bell instability generates a strong magnetic field perpendicular to the shock normal, which then changes the local magnetic field angle (Crumley2019). Although the shock started subluminal, this effect can make it superluminal for the low-energy electrons that have a Larmor radius smaller than the half-wavelength of Bell modes. Thus, electron injection is suppressed. The filamentary structure of nonlinear Bell modes still allows a fraction of electrons to be injected, but a significant decrease in the injection rate is inevitable, due to the superluminality constraint mentioned above. Ions could still experience efficient injection because they have a much larger Larmor radius, even for the low-energy population at energies just above the thermal peak. High-energy electrons, which already had a Larmor radius larger than the Bell scale, can still be accelerated.
@@
Figure 8 shows the time evolution of the maximum particle energy $E_{\mathrm{max},s}=m_{s}c^{2}(\gamma_{\mathrm{max}}-1)$, where $\gamma_{\mathrm{max}}$ is the Lorentz factor at which $(\gamma-1)f(\gamma)$ becomes $10^{5}$ times smaller than the peak. Panel (a) shows the $E_{\mathrm{max,i}}(t)/m_{\mathrm{i}}c^{2}$ of ions. Ions are efficiently accelerated in all cases, but the scaling of the maximum Lorentz factor is different. The nonthermal tail is growing the fastest in the highest-$\sigma$ case and the slowest in the lowest-$\sigma$ case. In particular, we find a Bohm-like scaling $E_{\mathrm{max,i}}\propto t$ in the $\sigma=10^{-3}$ case (dark teal), whereas lower magnetization cases (dark green and turquoise) follow the $E_{\mathrm{max,i}}\propto t^{1/2}$ scaling. The latter is consistent with ultra-relativistic shock simulations by Sironi2013, who found that the maximum energy of particles accelerated in small-scale turbulence, such as the Weibel fields, scales as $E_{\mathrm{max}}\propto t^{1/2}$. In contrast, the Bohm scaling $E_{\mathrm{max}}\propto t$ is expected for Bell-dominated shocks (Drury1983; Gargate2012; Stockem2012).
@@
Figure 9: Magnetization dependence of the downstream kinetic energy density. Panels (a) and (b) show the total energy density (Equation (6)) and the nonthermal energy density (Equation (7)), respectively, in the downstream region. The yellow and blue lines correspond to ions and electrons, respectively.
Figure 9 (a) shows the total energy density of each particle species. The three low-magnetization cases $(\sigma\lesssim 10^{-3.5})$ show $\mathcal{E}_{\mathrm{i}}\sim 0.6$ and $0.3\lesssim\mathcal{E}_{\mathrm{e}}\lesssim 0.4$, respectively. Electron heating to near equipartition with ions is consistent with the result of unmagnetized shocks in both ultra-relativistic (Spitkovsky2008a; Vanthieghem2022) and non-relativistic (Vanthieghem2024) regimes. The Bell-dominated $\sigma=10^{-3}$ case exhibits a noticeably different energy partition $\mathcal{E}_{\mathrm{i}}\sim 0.8$ and $\mathcal{E}_{\mathrm{e}}\sim 0.2$, in other words, weaker electron heating.
The energy density of the nonthermal component is shown in Figure 9(b). For all four magnetizations, the ion nonthermal energy is $0.1\lesssim\varepsilon_{\mathrm{i}}\lesssim 0.2$, consistent with previous simulations in a wide range of $\Gamma_{\mathrm{sh}}$ and $\sigma$ (e.g., Sironi2013; Caprioli2014a; Crumley2019). Note that the precise values of $\varepsilon_{\mathrm{i}}$ and $\varepsilon_{\mathrm{e}}$ depend on our choice of $\gamma_{\mathrm{inj}}$. Therefore, we shall mostly focus on the dependence on $\sigma$, which is rather weak for the ion species.
@@
The values for the three lower-$\sigma$ cases $(\sigma=10^{-4.5},10^{-4}$, and $10^{-3.5})$ are all around $m_{\mathrm{i}}/m_{\mathrm{e}}]_{\mathrm{eff}}\sim 5$, while $m_{\mathrm{i}}/m_{\mathrm{e}}]_{\mathrm{eff}}\sim 10$ for the $\sigma=10^{-3}$ case. This implies that our simulation mass ratio $m_{\mathrm{i}}/m_{\mathrm{e}}=100\gg[m_{\mathrm{i}}/m_{\mathrm{e}}]_{\mathrm{eff}}$ is sufficiently large to capture, both qualitatively and quantitatively, the key aspects of ion and electron energy partitioning. In other words, due to efficient heating to ultra-relativistic temperatures, electrons lose memory of the initial mass difference. Previous studies have confirmed this conclusion with mass ratio surveys for ultra-relativistic and trans-relativistic shocks (Spitkovsky2008a; Sironi2013; Crumley2019).
@@
4 Summary and Discussion
4.1 Summary of the Results
Here, we summarize our main results. Let us first summarize the three main results discussed in Section 3. Recall that our fixed parameters are: the shock Lorentz factor $\Gamma_{\mathrm{sh}}\sim 2$ and the angle $\theta_{B}=20^{\circ}$ between the shock normal and the ambient mean magnetic field, defined in the upstream frame. We have investigated the magnetization dependence in the range from $\sigma=10^{-4.5}$ to $\sigma=10^{-3}$. Our main findings are as follows:
@@ 4.2 Maximum Ion Energy
which makes them promising candidates for UHECR production (Cerutti2023; Globus2025). For a microquasar like SS 433,
@@
so the recent detection of $\sim 100\,\mathrm{TeV}$ gamma rays (Cao2025) could have a hadronic origin. Bell-dominated trans-relativistic shocks may also be relevant to ion acceleration and gamma ray emission in ultrafast outflows (UFOs) from AGN (Ajello2021; Nishiura2025).
For particle acceleration in these relatively highly magnetized shocks, it is essential that the shock velocity is trans-relativistic, instead of ultra-relativistic. As we have discussed in Section 1 and 2, ultra-relativistic shocks are inevitably superluminal. For any finite magnetization, the maximum energy in ultra-relativistic shocks is limited by the magnetization as $\gamma_{\mathrm{max}}\propto\sigma^{-1/4}$ (Sironi2013; Plotnikov2018). Note that, however, Reville2014 proposed a stronger scaling $(\gamma_{\mathrm{max}}\propto\sigma^{-1/2})$.
4.3 Synchrotron Emission
Here, we discuss thermal and nonthermal synchrotron emission from trans-relativistic astrophysical shocks. Let us start with the nonthermal emission from late-time GRB afterglows. Previous studies show that ultra-relativistic low-magnetization shocks can efficiently generate nonthermal electrons (Spitkovsky2008a; Sironi2013; Groselj2024). Observations of GW170817 indicate that a significant amount of nonthermal electron energy is required to model the late-time trans-relativistic stage of the GRB external shock (Margutti2018; Hajela2019; Hajela2022).
However, the PIC simulations by Crumley2019 imply that the electron nonthermal energy fraction is small, i.e., $\varepsilon_{\mathrm{e}}\ll 0.1$. We have shown that this only holds at relatively high ambient magnetizations $(\sigma\gtrsim 10^{-3})$, which are considerably higher than those expected in GRBs. Instead, a typical interstellar-like medium has a low magnetization:
@@ 4.3 Synchrotron Emission
Recent studies on FBOTs, such as AT2018cow (Ho2019; Margutti2019), reveal that synchrotron emission in those systems could instead be dominated by relativistic thermal electrons. Our results show substantial electron heating at all magnetizations. Furthermore, we have obtained microphysical parameters $\varepsilon_{B},\mathcal{E}_{\mathrm{e}},$ and $\varepsilon_{\mathrm{e}}$, in the trans-relativistic, weakly magnetized regime. These quantities can be used, for example, in models that infer shock velocity and ambient density from observed light curves (Margalit2021; Margalit2024).
4.4 Deceleration Signature
@@
In this paper, we have studied shocks with Lorentz factor $\Gamma_{\mathrm{sh}}\sim 2$. Here, we shall discuss the physics in different shock velocity regimes and how that can be used to extract information from decelerating shocks. Let us start with the ultra-relativistic limit $(\Gamma_{\mathrm{sh}}\gg 1)$. In this case, the shock will be quasi-perpendicular, regardless of the upstream $\theta_{B}$. This regime has been investigated extensively both for electron-positron and electron-ion plasmas (Kato2007; Kato2008; Spitkovsky2008a; Spitkovsky2008b; Sironi2013; Groselj2022). The shock is Weibel-dominated, and thus the maximum particle energy has a $E_{\mathrm{max}}\propto t^{1/2}$ scaling. It is worth pointing out that moderately magnetized $10^{-4}\lesssim\sigma\lesssim 10^{-3}$ electron-ion shocks are still relatively under-explored. In this magnetization range, the Weibel instability may be modified by the ambient field, affecting the electron physics (Jikei2024a). In non-relativistic shocks, this can trigger magnetic reconnection near the shock front (Matsumoto2015; Bohdan2021). Similar arguments may also apply to relativistic shocks, especially at relatively low Lorentz factors $\Gamma_{\mathrm{sh}}\lesssim 10$.
@@
We now move to shock Lorentz factors smaller than $\Gamma_{\mathrm{sh}}\sim 2$. In the case of non-relativistic shocks, we expect a quasi-parallel shock to be Bell-dominated even at very low magnetization, or equivalently, at very high Alfvénic Mach number $M_{\mathrm{A}}$. Hybrid simulations show efficient magnetic amplification by the Bell instability (Caprioli2014b). Intense heating of the incoming electrons would be required for efficient magnetic field generation by the Weibel instability (Lyubarsky2006; Jikei2024b). Heating of electrons to a level that significantly reduces the effective mass ratio $m_{\mathrm{i}}/m_{\mathrm{e}}]_{\mathrm{eff}}$ is not possible in the non-relativistic regime. Although the suppression of electron injection by the superluminality constraint does not apply in the non-relativistic limit, the nonthermal electron energy fraction is still small $(\varepsilon_{\mathrm{e}}\ll 0.1)$ in non-relativistic Bell-dominated shocks (Park2015).
@@
These considerations lead us to the speculation that almost every relativistic shock wave starts as a Weibel-dominated shock and eventually transitions to a Bell-dominated state after deceleration to non-relativistic speeds. Since Weibel-dominated shocks convert significant energy into nonthermal electrons $(\varepsilon_{\mathrm{e}}\sim 0.1)$, while $\varepsilon_{\mathrm{e}}$ is much smaller for Bell-dominated shocks, we argue that the nonthermal luminosity should abruptly decrease when the shock decelerates to non-relativistic speeds. For example, the observation of relatively bright nonthermal emission for GW170817 for more than 1000 days (Hajela2019; Hajela2022) is consistent with the common assumption that the upstream environment is low-$\sigma$ and that the shock speed is still trans-relativistic. As a direct follow-up to this work, we will characterize the transition (in shock speed) from a Weibel-dominated to a Bell-dominated shock.

### v2

@@ III.2 Cosmic Ray Driven Instabilities
This corresponds to the simulation frame magnetic field energy density divided by the upstream kinetic energy density. Note that this is a frame-dependent quantity. The strength of the near upstream Weibel field seen at $(x-x_{\mathrm{sh}})/d_{\mathrm{i}}=[200,500]$, in the dark green and turquoise curves is $\varepsilon_{B}\sim 2\times 10^{-4}$. The shock transition region of the $\sigma=10^{-4.5}$ case, as seen from the $\varepsilon_{B}$ profile, is slightly broader, compared to the $\sigma=10^{-4}$ case. This is due to the larger-scale structures generated during the strongly nonlinear stage of the Weibel shock evolution (see Appendix A and Grošelj et al. (2024)). For the $\sigma=10^{-3.5}$ case in the orange curve, we see some spikes peaking at $\varepsilon_{B}\sim 10^{-3}$ at $(x-x_{\mathrm{sh}})/d_{\mathrm{i}}=[200,300]$, which is what we observed in Figure 3(c). The upstream magnetic field of the $\sigma=10^{-3}$ case is much stronger at $\varepsilon_{B}\sim 10^{-2}$. The high-current Bell modes generated in the $\sigma=10^{-3.5}$ case saturate an order of magnitude below the low-current Bell modes of the $\sigma=10^{-3}$ case, but still remain locally larger than the Weibel modes, which dominate at lower magnetizations. In all cases, $\varepsilon_{B}$ approaches the far upstream value;
@@
In this Subsection, we analyzed the cosmic ray currents and the plasma instabilities driven by cosmic rays in the upstream region. By considering the Weibel instability and both low-current and high-current Bell instability, we were able to elucidate the $\sigma$-dependence on the shock structure seen in Subsection III.1. For additional details regarding the saturation of the high-current Bell modes, see Appendix B.
III.3 Particle Acceleration
@@
For electrons, the time evolution of the nonthermal spectra has a clear dependence on $\sigma$. For the low-$\sigma$ cases (panels (f,h)), the nonthermal tail has a relatively steady normalization, and its upper cutoff keeps growing. On the other hand, for the Bell-dominated case with $\sigma=10^{-3}$ (panel (b)), the normalization of the tail gets suppressed after $\omega_{\mathrm{pi}}t\sim 4000$. This is because the Bell instability generates a strong magnetic field perpendicular to the shock normal, which then changes the local magnetic field angle (Crumley et al., 2019). Although the shock started subluminal, this effect can make it superluminal for the low-energy electrons that have a Larmor radius smaller than the half-wavelength of Bell modes. Thus, electron injection is suppressed. The filamentary structure of nonlinear Bell modes still allows a fraction of electrons to be injected, but a significant decrease in the injection rate is inevitable, due to the superluminality constraint mentioned above. Ions could still experience efficient injection because they have a much larger Larmor radius, even for the low-energy population at energies just above the thermal peak. High-energy electrons, which already had a Larmor radius larger than the Bell scale, can still be accelerated.
@@
Figure 8 shows the time evolution of the maximum particle energy $E_{\mathrm{max},s}=m_{s}c^{2}(\gamma_{\mathrm{max}}-1)$, where $\gamma_{\mathrm{max}}$ is the Lorentz factor at which $(\gamma-1)f(\gamma)$ becomes $10^{5}$ times smaller than the peak. Panel (a) shows the $E_{\mathrm{max,i}}(t)/m_{\mathrm{i}}c^{2}$ of ions. Ions are efficiently accelerated in all cases, but the scaling of the maximum Lorentz factor is different. The nonthermal tail is growing the fastest in the highest-$\sigma$ case and the slowest in the lowest-$\sigma$ case. In particular, we find a Bohm-like scaling $E_{\mathrm{max,i}}\propto t$ in the $\sigma=10^{-3}$ case (dark teal), whereas lower magnetization cases (dark green and turquoise) follow the $E_{\mathrm{max,i}}\propto t^{1/2}$ scaling. The latter is consistent with ultra-relativistic shock simulations by Sironi et al. (2013), who found that the maximum energy of particles accelerated in small-scale turbulence, such as the Weibel fields, scales as $E_{\mathrm{max}}\propto t^{1/2}$. In contrast, the Bohm scaling $E_{\mathrm{max}}\propto t$ is expected for Bell-dominated shocks (Drury, 1983; Gargaté and Spitkovsky, 2012; Stockem et al., 2012).
@@
Figure 9: Magnetization dependence of the downstream kinetic energy partition and the fraction of nonthermal particles. Panel (a) shows the total energy density (Equation (6)), and panel (b) shows the nonthermal energy density (Equation (7), solid lines) and nonthermal number fraction (Equation (8), dotted lines), respectively, in the downstream region. The yellow and blue lines correspond to ions and electrons, respectively.
Figure 9 (a) shows the total energy density of each particle species. The three low-magnetization cases $(\sigma\lesssim 10^{-3.5})$ show $\mathcal{E}_{\mathrm{i}}\sim 0.6$ and $0.3\lesssim\mathcal{E}_{\mathrm{e}}\lesssim 0.4$, respectively. Electron heating to near equipartition with ions is consistent with the result of unmagnetized shocks in both ultra-relativistic (Spitkovsky, 2008a; Vanthieghem et al., 2022) and non-relativistic (Vanthieghem et al., 2024) regimes. The Bell-dominated $\sigma=10^{-3}$ case exhibits a noticeably different energy partition $\mathcal{E}_{\mathrm{i}}\sim 0.8$ and $\mathcal{E}_{\mathrm{e}}\sim 0.2$, in other words, weaker electron heating.
The energy density of the nonthermal component is shown in Figure 9(b) solid lines. For all four magnetizations, the ion nonthermal energy is $0.1\lesssim\varepsilon_{\mathrm{i}}\lesssim 0.2$, consistent with previous simulations in a wide range of $\Gamma_{\mathrm{sh}}$ and $\sigma$ (e.g., Sironi et al., 2013; Caprioli and Spitkovsky, 2014a; Crumley et al., 2019). Note that the precise values of $\varepsilon_{\mathrm{i}}$ and $\varepsilon_{\mathrm{e}}$ depend on our choice of $\gamma_{\mathrm{inj}}$. Therefore, we shall mostly focus on the dependence on $\sigma$, which is rather weak for the ion species.
@@
We define the downstream number fraction of the nonthermal particles of species $s$ as:
$\tilde{n}_{s}=\frac{\int_{\gamma_{\mathrm{inj}}}^{\infty}f_{s}d\gamma}{\int_{1}^{\infty}f_{s}d\gamma}.$ (8)
The dotted lines in Figure 9(b) show $\tilde{n}_{s}$. The values are between $0.04$ and $0.12$ for both ions and electrons, and have a similar $\sigma$-dependence as the energy fractions $\varepsilon_{s}$.
@@
The values for the three lower-$\sigma$ cases $(\sigma=10^{-4.5},10^{-4}$, and $10^{-3.5})$ are all around $m_{\mathrm{i}}/m_{\mathrm{e}}]_{\mathrm{eff}}\sim 5$, while $m_{\mathrm{i}}/m_{\mathrm{e}}]_{\mathrm{eff}}\sim 10$ for the $\sigma=10^{-3}$ case. This implies that our simulation mass ratio $m_{\mathrm{i}}/m_{\mathrm{e}}=100\gg[m_{\mathrm{i}}/m_{\mathrm{e}}]_{\mathrm{eff}}$ is sufficiently large to capture, both qualitatively and quantitatively, the key aspects of ion and electron energy partitioning. In other words, due to efficient heating to ultra-relativistic temperatures, electrons lose memory of the initial mass difference. Previous studies have confirmed this conclusion with mass ratio surveys for ultra-relativistic and trans-relativistic shocks (Spitkovsky, 2008a; Sironi et al., 2013; Crumley et al., 2019).
@@
IV Summary and Discussion
IV.1 Summary of the Results
Here, we summarize our main results. Let us first summarize the three main results discussed in Section III. Recall that our fixed parameters are: the shock Lorentz factor $\Gamma_{\mathrm{sh}}\sim 2$ and the angle $\theta_{B}=20^{\circ}$ between the shock normal and the ambient mean magnetic field, defined in the upstream frame. We have investigated the magnetization dependence in the range from $\sigma=10^{-4.5}$ to $\sigma=10^{-3}$. Our main findings are as follows:
@@ IV.2 Maximum Ion Energy
which makes them promising candidates for UHECR production (Cerutti and Giacinti, 2023; Globus and Blandford, 2025). For a microquasar like SS 433,
@@
so the recent detection of $\sim 100\,\mathrm{TeV}$ gamma rays (Cao et al., 2025) could have a hadronic origin. Bell-dominated trans-relativistic shocks may also be relevant to ion acceleration and gamma ray emission in ultrafast outflows (UFOs) from AGN (Ajello et al., 2021; Nishiura and Inoue, 2026). Applications to the nonthermal emission from BL Lacertae objects have also been discussed (Crumley et al., 2019; Arbet-Engels et al., 2025).
For particle acceleration in these relatively highly magnetized shocks, it is essential that the shock velocity is trans-relativistic, instead of ultra-relativistic. As we have discussed in Section I and II, ultra-relativistic shocks are inevitably superluminal. For any finite magnetization, the maximum energy in ultra-relativistic shocks is limited by the magnetization as $\gamma_{\mathrm{max}}\propto\sigma^{-1/4}$ (Sironi et al., 2013; Plotnikov et al., 2018). Note that, however, Reville and Bell (2014) proposed a stronger scaling $(\gamma_{\mathrm{max}}\propto\sigma^{-1/2})$.
IV.3 Synchrotron Emission
Here, we discuss thermal and nonthermal synchrotron emission from trans-relativistic astrophysical shocks. Let us start with the nonthermal emission from late-time GRB afterglows. Previous studies show that ultra-relativistic low-magnetization shocks can efficiently generate nonthermal electrons (Spitkovsky, 2008a; Sironi et al., 2013; Grošelj et al., 2024). Observations of GW170817 indicate that a significant amount of nonthermal electron energy is required to model the late-time trans-relativistic stage of the GRB external shock (Margutti et al., 2018; Hajela et al., 2019, 2022).
However, the PIC simulations by Crumley et al. (2019) imply that the electron nonthermal energy fraction is small, i.e., $\varepsilon_{\mathrm{e}}\ll 0.1$. We have shown that this only holds at relatively high ambient magnetizations $(\sigma\gtrsim 10^{-3})$, which are considerably higher than those expected in GRBs. Instead, a typical interstellar-like medium has a low magnetization:
@@ IV.3 Synchrotron Emission
Recent studies on FBOTs, such as AT2018cow (Ho et al., 2019; Margutti et al., 2019), reveal that synchrotron emission in those systems could instead be dominated by relativistic thermal electrons. Our results show substantial electron heating at all magnetizations. Furthermore, we have obtained microphysical parameters $\varepsilon_{B},\mathcal{E}_{\mathrm{e}},$ and $\varepsilon_{\mathrm{e}}$, in the trans-relativistic, weakly magnetized regime. These quantities can be used, for example, in models that infer shock velocity and ambient density from observed light curves (Margalit and Quataert, 2021, 2024).
IV.4 Deceleration Signature
In this paper, we have studied shocks with Lorentz factor $\Gamma_{\mathrm{sh}}\sim 2$. Here, we shall discuss the physics in different shock velocity regimes and how that can be used to extract information from decelerating shocks. Here, we refer to the deceleration of astrophysical shocks as they progressively slow down while propagating into their surrounding medium, on dynamical timescales (therefore, much longer than the plasma scales of our simulations). Time evolution of macroscopic structures, during ultra-relativistic and non-relativistic stages, can be described by the Blandford-McKee solution (Blandford and McKee, 1976) and the classical Taylor-von Neumann-Sedov solution (e.g., Landau and Lifshitz, 1987), respectively.
@@ IV.4 Deceleration Signature
Let us start with the ultra-relativistic limit $(\Gamma_{\mathrm{sh}}\gg 1)$. In this case, the shock will be quasi-perpendicular, regardless of the upstream $\theta_{B}$. This regime has been investigated extensively both for electron-positron and electron-ion plasmas (Kato, 2007; Kato and Takabe, 2008; Spitkovsky, 2008a, b; Sironi et al., 2013; Grošelj et al., 2022). The shock is Weibel-dominated, and thus the maximum particle energy has a $E_{\mathrm{max}}\propto t^{1/2}$ scaling. It is worth pointing out that moderately magnetized $10^{-4}\lesssim\sigma\lesssim 10^{-3}$ electron-ion shocks are still relatively under-explored. In this magnetization range, the Weibel instability may be modified by the ambient field, affecting the electron physics (Jikei et al., 2024). In non-relativistic shocks, this can trigger magnetic reconnection near the shock front (Matsumoto et al., 2015; Bohdan et al., 2021). Similar arguments may also apply to relativistic shocks, especially at relatively low Lorentz factors $\Gamma_{\mathrm{sh}}\lesssim 10$.
@@
We now move to shock Lorentz factors smaller than $\Gamma_{\mathrm{sh}}\sim 2$. In the case of non-relativistic shocks, we expect a quasi-parallel shock to be Bell-dominated even at very low magnetization, or equivalently, at very high Alfvénic Mach number $M_{\mathrm{A}}$. Hybrid simulations show efficient magnetic amplification by the Bell instability (Caprioli and Spitkovsky, 2014b). Intense heating of the incoming electrons would be required for efficient magnetic field generation by the Weibel instability (Lyubarsky and Eichler, 2006; Jikei and Amano, 2024). Heating of electrons to a level that significantly reduces the effective mass ratio $m_{\mathrm{i}}/m_{\mathrm{e}}]_{\mathrm{eff}}$ is not possible in the non-relativistic regime. Although the suppression of electron injection by the superluminality constraint does not apply in the non-relativistic limit, the nonthermal electron energy fraction is still small $(\varepsilon_{\mathrm{e}}\ll 0.1)$ in non-relativistic Bell-dominated shocks (Park et al., 2015).
@@
These considerations lead us to the speculation that almost every relativistic shock wave starts as a Weibel-dominated shock and eventually transitions to a Bell-dominated state after deceleration to non-relativistic speeds. Since Weibel-dominated shocks convert significant energy into nonthermal electrons $(\varepsilon_{\mathrm{e}}\sim 0.1)$, while $\varepsilon_{\mathrm{e}}$ is much smaller for Bell-dominated shocks, we argue that the nonthermal luminosity should abruptly decrease when the shock decelerates to non-relativistic speeds. For example, the observation of relatively bright nonthermal emission for GW170817 for more than 1000 days (Hajela et al., 2019, 2022) is consistent with the common assumption that the upstream environment is low-$\sigma$ and that the shock speed is still trans-relativistic. As a direct follow-up to this work, we will characterize the transition (in shock speed) from a Weibel-dominated to a Bell-dominated shock.